OpenStudy (anonymous):

For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept.

OpenStudy (anonymous):

To find y intercept, you want to see where the point is when x=0. To find where the x intercept is, you want to see where the point is when y=0

OpenStudy (anonymous):

@Denebel, but I do not have a Graph

OpenStudy (anonymous):

You don't need to graph, all you do is: y = -2x - 3 To find x intercept, plug 0 in y: 0=-2x-3.... Solve so you get x=_______ and that is your x-intercept. To find y, do the opposite: y=-2(0)-3 Get y=_______ and that is your y-intercept. On a side note, for linear equations, y=mx+b, you would know that b will be your y-intercept (see above for explanation).

OpenStudy (anonymous):

So the x-intercept would be \[-3\over 2\]

OpenStudy (anonymous):

And Y-intercept would be 0.5? @Denebel

OpenStudy (anonymous):

x-intercept is correct but y=-2(0)-3 y=0-3 y=-3

OpenStudy (anonymous):

Is that all I need for my ANswer @Denebel

OpenStudy (anonymous):

Yes those points will be your x and y intercepts

OpenStudy (anonymous):

Thank so much @Denebel